# An insight on some properties of high order nonstandard linear multistep methods

### v1

@@ 5.2 The SEIR system
$0.1$ $1.1531\cdot 10^{-1}$ $8.5437\cdot 10^{-2}$ $2.0912\cdot 10^{-1}$
$0.1\cdot 2^{-1}$ $5.5339\cdot 10^{-2}$ $1.0592$ $4.0419\cdot 10^{-2}$ $1.0798$ $1.0887\cdot 10^{-1}$ $0.9416$
$0.1\cdot 2^{-2}$ $2.6622\cdot 10^{-2}$ $1.0556$ $1.9456\cdot 10^{-2}$ $1.0548$ $5.3511\cdot 10^{-2}$ $1.0247$
$0.1\cdot 2^{-3}$ $1.2993\cdot 10^{-2}$ $1.0349$ $9.5200\cdot 10^{-3}$ $1.0312$ $2.6159\cdot 10^{-2}$ $1.0325$
$0.1\cdot 2^{-4}$ $6.4107\cdot 10^{-3}$ $1.0192$ $4.7058\cdot 10^{-3}$ $1.0165$ $1.2878\cdot 10^{-2}$ $1.0224$
$0.1\cdot 2^{-5}$ $3.1830\cdot 10^{-3}$ $1.0101$ $2.3391\cdot 10^{-3}$ $1.0085$ $6.3822\cdot 10^{-3}$ $1.0128$
$0.1\cdot 2^{-6}$ $1.5859\cdot 10^{-3}$ $1.0051$ $1.1661\cdot 10^{-3}$ $1.0043$ $3.1760\cdot 10^{-3}$ $1.0069$
$0.1\cdot 2^{-7}$ $7.9150\cdot 10^{-4}$ $1.0026$ $5.8217\cdot 10^{-4}$ $1.0022$ $1.5841\cdot 10^{-3}$ $1.0035$
$0.1\cdot 2^{-8}$ $3.9539\cdot 10^{-4}$ $1.0013$ $2.9087\cdot 10^{-4}$ $1.0011$ $7.9106\cdot 10^{-3}$ $1.0018$
$0.1\cdot 2^{-9}$ $1.9761\cdot 10^{-4}$ $1.0007$ $1.4538\cdot 10^{-4}$ $1.0005$ $3.9528\cdot 10^{-4}$ $1.0009$
@@
$0.1$ $8.6047\cdot 10^{-2}$ $3.9627\cdot 10^{-2}$ $5.6965\cdot 10^{-2}$
$0.1\cdot 2^{-1}$ $2.4040\cdot 10^{-2}$ $1.8397$ $1.0037\cdot 10^{-2}$ $1.9811$ $1.4924\cdot 10^{-2}$ $1.9324$
$0.1\cdot 2^{-2}$ $6.1921\cdot 10^{-3}$ $1.9570$ $2.5270\cdot 10^{-3}$ $1.9898$ $3.7831\cdot 10^{-3}$ $1.9800$
$0.1\cdot 2^{-3}$ $1.5642\cdot 10^{-3}$ $1.9850$ $6.3499\cdot 10^{-4}$ $1.9926$ $9.5204\cdot 10^{-4}$ $1.9905$
$0.1\cdot 2^{-4}$ $3.9275\cdot 10^{-4}$ $1.9937$ $1.5924\cdot 10^{-4}$ $1.9955$ $2.3884\cdot 10^{-4}$ $1.9950$
$0.1\cdot 2^{-5}$ $9.8386\cdot 10^{-5}$ $1.9971$ $3.9878\cdot 10^{-5}$ $1.9975$ $5.9816\cdot 10^{-5}$ $1.9974$
$0.1\cdot 2^{-6}$ $2.4620\cdot 10^{-5}$ $1.9986$ $9.9785\cdot 10^{-6}$ $1.9987$ $1.4968\cdot 10^{-5}$ $1.9987$
$0.1\cdot 2^{-7}$ $6.1580\cdot 10^{-6}$ $1.9993$ $2.4958\cdot 10^{-6}$ $1.9993$ $3.7436\cdot 10^{-6}$ $1.9993$
$0.1\cdot 2^{-8}$ $1.5399\cdot 10^{-6}$ $1.9997$ $6.2409\cdot 10^{-7}$ $1.9997$ $9.3613\cdot 10^{-7}$ $1.9997$
$0.1\cdot 2^{-9}$ $3.8502\cdot 10^{-7}$ $1.9998$ $1.5605\cdot 10^{-7}$ $1.9997$ $2.3407\cdot 10^{-7}$ $1.9998$
@@
$0.1$ $2.2574\cdot 10^{-2}$ $1.0324\cdot 10^{-2}$
$0.1\cdot 2^{-1}$ $3.0024\cdot 10^{-3}$ $2.9105$ $6.8634\cdot 10^{-4}$ $3.9110$
$0.1\cdot 2^{-2}$ $3.8218\cdot 10^{-4}$ $2.9738$ $4.3684\cdot 10^{-5}$ $3.9737$
$0.1\cdot 2^{-3}$ $4.8138\cdot 10^{-5}$ $2.9890$ $2.7538\cdot 10^{-6}$ $3.9876$
$0.1\cdot 2^{-4}$ $6.0390\cdot 10^{-6}$ $2.9948$ $1.7287\cdot 10^{-7}$ $3.9937$
$0.1\cdot 2^{-5}$ $7.5622\cdot 10^{-7}$ $2.9974$ $1.0829\cdot 10^{-8}$ $3.9967$
$0.1\cdot 2^{-6}$ $9.4613\cdot 10^{-8}$ $2.9987$ $6.7923\cdot 10^{-10}$ $3.9949$
$0.1\cdot 2^{-7}$ $1.1835\cdot 10^{-8}$ $2.9990$ $4.5636\cdot 10^{-11}$ $3.8957$
$0.1\cdot 2^{-8}$ $1.4857\cdot 10^{-9}$ $2.9939$ $9.0434\cdot 10^{-12}$ $2.3352$
$0.1\cdot 2^{-9}$ $1.9758\cdot 10^{-10}$ $2.9106$ $1.2814\cdot 10^{-11}$
Table 3: The errors and the corresponding orders of the different methods using different $\varphi$ functions.
[…]
Figure 5: The errors of the different multistep methods for different values of $\Delta t$.
NSSPMS(4,2) NSSPMS(4,3) NSSPMS(6,4)
@@
$0.05$ $3.7854\cdot 10^{-4}$ $4.9316\cdot 10^{-4}$ $6.1934\cdot 10^{-4}$
$0.05\cdot 2^{-1}$ $1.0095\cdot 10^{-4}$ $1.9068$ $9.1717\cdot 10^{-6}$ $2.4267$ $4.2076\cdot 10^{-5}$ $3.8797$
$0.05\cdot 2^{-2}$ $2.6162\cdot 10^{-5}$ $1.9481$ $1.4250\cdot 10^{-6}$ $2.6862$ $2.7473\cdot 10^{-6}$ $3.9369$
$0.05\cdot 2^{-3}$ $6.6660\cdot 10^{-6}$ $1.9726$ $1.9672\cdot 10^{-7}$ $2.8568$ $1.7543\cdot 10^{-7}$ $3.9690$
$0.05\cdot 2^{-4}$ $1.6829\cdot 10^{-6}$ $1.9859$ $2.5792\cdot 10^{-8}$ $2.9312$ $1.1084\cdot 10^{-8}$ $3.9845$
$0.05\cdot 2^{-5}$ $4.2280\cdot 10^{-7}$ $1.9928$ $3.3003\cdot 10^{-9}$ $2.9662$ $6.9645\cdot 10^{-10}$ $3.9923$
$0.05\cdot 2^{-6}$ $1.0597\cdot 10^{-7}$ $1.9964$ $4.1735\cdot 10^{-10}$ $2.9833$ $4.3562\cdot 10^{-11}$ $3.9989$
$0.05\cdot 2^{-7}$ $2.6525\cdot 10^{-8}$ $1.9982$ $5.2470\cdot 10^{-11}$ $2.9917$ $2.5479\cdot 10^{-12}$ $4.0957$
$0.05\cdot 2^{-8}$ $6.6354\cdot 10^{-9}$ $1.9991$ $6.5846\cdot 10^{-12}$ $2.9943$ $1.1435\cdot 10^{-12}$ $1.1558$

### v2

@@ Appendix B. Tables of errors and orders
$2\cdot 10^{-4}$ $4.6308\cdot 10^{1}$ - $3.7984\cdot 10^{1}$ - $6.6018\cdot 10^{1}$ -
$2\cdot 10^{-4}\cdot 2^{-1}$ $3.5903\cdot 10^{1}$ $0.3672$ $2.7737\cdot 10^{1}$ $0.4536$ $6.6018\cdot 10^{1}$ $0.1555$
$2\cdot 10^{-4}\cdot 2^{-2}$ $1.9406\cdot 10^{1}$ $0.8876$ $1.4576\cdot 10^{1}$ $0.9283$ $3.5249\cdot 10^{1}$ $0.7498$
$2\cdot 10^{-4}\cdot 2^{-3}$ $1.0111\cdot 10^{1}$ $0.9405$ $7.5087\cdot 10^{0}$ $0.9570$ $1.9283\cdot 10^{1}$ $0.8702$
$2\cdot 10^{-4}\cdot 2^{-4}$ $5.1489\cdot 10^{0}$ $0.9736$ $3.8049\cdot 10^{0}$ $0.9807$ $1.0059\cdot 10^{0}$ $0.9389$
$2\cdot 10^{-4}\cdot 2^{-5}$ $2.5969\cdot 10^{0}$ $0.9875$ $1.9147\cdot 10^{0}$ $0.9907$ $5.1337\cdot 10^{0}$ $0.9704$
$2\cdot 10^{-4}\cdot 2^{-6}$ $1.3040\cdot 10^{0}$ $0.9939$ $9.6039\cdot 10^{-1}$ $0.9955$ $2.5928\cdot 10^{0}$ $0.9855$
$2\cdot 10^{-4}\cdot 2^{-7}$ $6.5334\cdot 10^{-1}$ $0.9970$ $4.8094\cdot 10^{-1}$ $0.9977$ $1.3029\cdot 10^{0}$ $0.9928$
$2\cdot 10^{-4}\cdot 2^{-8}$ $3.2701\cdot 10^{-1}$ $0.9985$ $2.4066\cdot 10^{-1}$ $0.9989$ $6.5307\cdot 10^{-1}$ $0.9964$
$2\cdot 10^{-4}\cdot 2^{-9}$ $1.6359\cdot 10^{-1}$ $0.9993$ $1.2038\cdot 10^{-1}$ $0.9994$ $3.2694\cdot 10^{-1}$ $0.9982$
@@
$2\cdot 10^{-4}$ $4.7539\cdot 10^{1}$ - $3.1548\cdot 10^{1}$ - $3.8509\cdot 10^{1}$ -
$2\cdot 10^{-4}\cdot 2^{-1}$ $2.7909\cdot 10^{1}$ $0.7684$ $3964\cdot 10^{1} 1.1$7591.1759 $1.9400\cdot 10^{1}$ $0.9891$
$2\cdot 10^{-4}\cdot 2^{-2}$ $8.9750\cdot 10^{0}$ $1.6368$ $3.8755\cdot 10^{0}$ $1.8492$ $5.6923\cdot 10^{0}$ $1.7690$
$2\cdot 10^{-4}\cdot 2^{-3}$ $2.4687\cdot 10^{0}$ $1.8622$ $1.0173\cdot 10^{0}$ $1.9296$ $1.5179\cdot 10^{0}$ $1.9070$
$2\cdot 10^{-4}\cdot 2^{-4}$ $6.3821\cdot 10^{-1}$ $1.9517$ $2.5975\cdot 10^{-1}$ $1.9696$ $3.8910\cdot 10^{-1}$ $1.9638$
$2\cdot 10^{-4}\cdot 2^{-5}$ $1.6165\cdot 10^{-1}$ $1.9812$ $6.5583\cdot 10^{-2}$ $1.9857$ $9.8342\cdot 10^{-2}$ $1.9843$
$2\cdot 10^{-4}\cdot 2^{-6}$ $4.0639\cdot 10^{-2}$ $1.9919$ $1.6475\cdot 10^{-2}$ $1.9931$ $2.4710\cdot 10^{-2}$ $1.9927$
$2\cdot 10^{-4}\cdot 2^{-7}$ $1.0186\cdot 10^{-2}$ $1.9963$ $4.1284\cdot 10^{-3}$ $1.9966$ $6.1925\cdot 10^{-3}$ $1.9965$
$2\cdot 10^{-4}\cdot 2^{-8}$ $2.5496\cdot 10^{-3}$ $1.9982$ $1.0333\cdot 10^{-3}$ $1.9983$ $1.5500\cdot 10^{-3}$ $1.9983$
$2\cdot 10^{-4}\cdot 2^{-9}$ $6.3778\cdot 10^{-4}$ $1.9991$ $2.5848\cdot 10^{-4}$ $1.9992$ $3.8772\cdot 10^{-4}$ $1.9992$
@@
$2\cdot 10^{-4}$ $2.9078\cdot 10^{1}$ - $2.4506\cdot 10^{1}$ -
$2\cdot 10^{-4}\cdot 2^{-1}$ $8.2661\cdot 10^{0}$ $1.8146$ $3.9029\cdot 10^{0}$ $2.6505$
$2\cdot 10^{-4}\cdot 2^{-2}$ $1.1812\cdot 10^{0}$ $2.8070$ $2.7194\cdot 10^{-1}$ $3.8432$
$2\cdot 10^{-4}\cdot 2^{-3}$ $1.5476\cdot 10^{-1}$ $2.9321$ $1.7704\cdot 10^{-2}$ $3.9411$
$2\cdot 10^{-4}\cdot 2^{-4}$ $1.9721\cdot 10^{-2}$ $2.9722$ $1.1274\cdot 10^{-3}$ $3.9730$
$2\cdot 10^{-4}\cdot 2^{-5}$ $2.4881\cdot 10^{-3}$ $2.9867$ $7.1124\cdot 10^{-5}$ $3.9865$
$2\cdot 10^{-4}\cdot 2^{-6}$ $3.1244\cdot 10^{-4}$ $2.9934$ $4.4661\cdot 10^{-6}$ $3.9932$
$2\cdot 10^{-4}\cdot 2^{-7}$ $3.9144\cdot 10^{-5}$ $2.9967$ $2.7957\cdot 10^{-7}$ $3.9977$
$2\cdot 10^{-4}\cdot 2^{-8}$ $4.8982\cdot 10^{-6}$ $2.9985$ $1.7056\cdot 10^{-8}$ $4.0349$
$2\cdot 10^{-4}\cdot 2^{-9}$ $6.1178\cdot 10^{-7}$ $3.0012$ $1.8974\cdot 10^{-10}$ $6.4901$
Table 4: The errors and the corresponding orders of method NSSPMS(6,4) using different $\varphi$ functions while solving the logistic equation with $c=500$, plotted in Figure 5.
NSSPMS(4,2) NSSPMS(4,3) with $\varphi_{8}$ NSSPMS(4,3) with $\varphi_{7}$
@@
$2\cdot 10^{-4}$ $2.9338\cdot 10^{0}$ - $2.9774\cdot 10^{0}$ - $7.1664\cdot 10^{0}$ -
$2\cdot 10^{-4}\cdot 2^{-1}$ $8.2578\cdot 10^{-1}$ $1.8289$ $1.5806\cdot 10^{-1}$ $4.2355$ $1.0250\cdot 10^{0}$ $2.8057$
$\cdot 10^{-4}\cdot 2^{-2} 2.2$$2313\cdot 10^{-1}$ $1.8879$ $3.4056\cdot 10^{-3}$ $5.5364$ $1.3489\cdot 10^{-1}$ $2.9257$
$2\cdot 10^{-4}\cdot 2^{-3}$ $5.8223\cdot 10^{-2}$ $1.9382$ $6.7605\cdot 10^{-4}$ $2.3327$ $1.7226\cdot 10^{-2}$ $2.9692$
$2\cdot 10^{-4}\cdot 2^{-4}$ $1.4886\cdot 10^{-2}$ $1.9677$ $1.5597\cdot 10^{-4}$ $2.1159$ $2.1757\cdot 10^{-3}$ $2.9850$
$2\cdot 10^{-4}\cdot 2^{-5}$ $3.7643\cdot 10^{-3}$ $1.9835$ $2.4047\cdot 10^{-5}$ $2.6973$ $2.7337\cdot 10^{-4}$ $2.9925$
$2\cdot 10^{-4}\cdot 2^{-6}$ $9.4654\cdot 10^{-4}$ $1.9916$ $3.2931\cdot 10^{-6}$ $2.8684$ $3.4260\cdot 10^{-5}$ $2.9963$
$2\cdot 10^{-4}\cdot 2^{-7}$ $2.3733\cdot 10^{-4}$ $1.9958$ $4.2969\cdot 10^{-7}$ $2.9381$ $4.2880\cdot 10^{-6}$ $2.9981$
$2\cdot 10^{-4}\cdot 2^{-8}$ $5.9418\cdot 10^{-5}$ $1.9979$ $5.4860\cdot 10^{-8}$ $2.9694$ $5.3633\cdot 10^{-7}$ $2.9991$
NSSPRK(2,2) NSSPRK(3,3) with $\varphi_{8}$ NSSPRK(3,3) with $\varphi_{7}$
$\Delta t$ errors orders errors orders errors orders
$2\cdot 10^{-4}$ $9.1774\cdot 10^{-1}$ - $4.8634\cdot 10^{-3}$ - $3.0723\cdot 10^{-1}$ -
$2\cdot 10^{-4}\cdot 2^{-1}$ $2.0672\cdot 10^{-1}$ $2.1503$ $3.4867\cdot 10^{-3}$ $4.8010$ $3.9080\cdot 10^{-2}$ $2.9749$
$2\cdot 10^{-4}\cdot 2^{-2}$ $4.9515\cdot 10^{-2}$ $2.0618$ $6.1653\cdot 10^{-4}$ $2.4996$ $4.9217\cdot 10^{-3}$ $2.9892$
$2\cdot 10^{-4}\cdot 2^{-3}$ $1.2152\cdot 10^{-2}$ $2.0267$ $8.8392\cdot 10^{-5}$ $2.8022$ $6.1740\cdot 10^{-4}$ $2.9949$
$2\cdot 10^{-4}\cdot 2^{-4}$ $3.0122\cdot 10^{-3}$ $2.0122$ $1.1758\cdot 10^{-5}$ $2.9103$ $7.7309\cdot 10^{-5}$ $2.9975$
$2\cdot 10^{-4}\cdot 2^{-5}$ $7.5003\cdot 10^{-4}$ $2.0058$ $1.5141\cdot 10^{-6}$ $2.9571$ $9.6720\cdot 10^{-6}$ $2.9988$
$2\cdot 10^{-4}\cdot 2^{-6}$ $1.8714\cdot 10^{-4}$ $2.0028$ $1.9205\cdot 10^{-7}$ $2.9789$ $1.2095\cdot 10^{-6}$ $2.9994$
$2\cdot 10^{-4}\cdot 2^{-7}$ $4.6739\cdot 10^{-5}$ $2.0014$ $2.4199\cdot 10^{-8}$ $2.9884$ $1.5120\cdot 10^{-7}$ $2.9999$
$2\cdot 10^{-4}\cdot 2^{-8}$ $1.1679\cdot 10^{-5}$ $2.0007$ $3.0757\cdot 10^{-9}$ $2.9760$ $1.8859\cdot 10^{-8}$ $3.0031$
NSSPRK(6,4) NSSPRK(10,4)
$\Delta t$ errors orders errors orders
$2\cdot 10^{-4}$ $2.4506\cdot 10^{1}$ - $1.2566\cdot 10^{-4}$ -
$2\cdot 10^{-4}\cdot 2^{-1}$ $3.9029\cdot 10^{0}$ $2.6505$ $7.8481\cdot 10^{-6}$ $4.0010$
$2\cdot 10^{-4}\cdot 2^{-2}$ $2.7194\cdot 10^{-1}$ $3.8432$ $4.8991\cdot 10^{-7}$ $4.0017$
$2\cdot 10^{-4}\cdot 2^{-3}$ $1.7704\cdot 10^{-2}$ $3.9411$ $3.0594\cdot 10^{-8}$ $4.0012$
$2\cdot 10^{-4}\cdot 2^{-4}$ $1.1274\cdot 10^{-3}$ $3.9730$ $1.9088\cdot 10^{-9}$ $4.0025$
$2\cdot 10^{-4}\cdot 2^{-5}$ $7.1124\cdot 10^{-5}$ $3.9865$ $1.1539\cdot 10^{-10}$ $4.0481$
$2\cdot 10^{-4}\cdot 2^{-6}$ $4.4661\cdot 10^{-6}$ $3.9932$ $2.2737\cdot 10^{-13}$ $8.9873$
$2\cdot 10^{-4}\cdot 2^{-7}$ $2.7957\cdot 10^{-7}$ $3.9977$ $1.6257\cdot 10^{-11}$ -
$2\cdot 10^{-4}\cdot 2^{-8}$ $1.7056\cdot 10^{-8}$ $4.0349$ $2.6034\cdot 10^{-11}$ -